# Good News for Polynomial Root-finding

### v35

@@ 10.4 The distance to a closest zero of $p$ and fast inclusion tests
For a complex $c$ and positive integers $j$ and $h$, Eqn. (1.13) and Observation 10.2 combined imply the following relationships independent of $p_{d}$:
@@ 11.1 Overview
Map (1.4) and Cor. 1.3 together reduce an e/i test for any disc to the test for the unit disc $D=D(0,1)$, and we devise soft e/i tests for that disc assuming that #$\#(D(0,2))\leq m$ for a fixed $m$, $1\leq m\leq d$ (cf. claim (iv) of Cor. 1.2). Our Alg. 11.1 of Sec. 11.2 and Alg. 11.2 of Sec. 11.3 are deterministic specifications of Alg. 1.1, which is always correct when it claims inclusion, by virtue of Lemma 1.1. We prove that both of these specifications are also correct when they output exclusion. Namely, Algs. 11.1 and 11.2 evaluate ${\rm NIR}(x)$ at $hq$ equally spaced points of the unit circle $C(0,1)$ for $h=O(\log(d))$ and correctly claim exclusion if at all these points $|{\rm NIR}(x)|\leq\frac{1}{2\sqrt{q}}$ for $q>d$ or if $|{\rm NIR}(x)|\leq\frac{1}{3\sqrt{q}}$ for $q>m$, respectively. Our Example 11.1 shows that we cannot devise such deterministic algorithms for $q$ decreased below $m$, but in Sec. 11.4 we devise Las Vegas randomized e/i tests that evaluate polynomial $p(x)$ at $O(\log(d))$ points $x$ overall under the Random Root Model.
@@ 11.2 Basic deterministic e/i test
INPUT: a polynomial $p(x)$ of a degree $d$ and $\sigma$, $1<\sigma<\sqrt{2}$.
@@
Alg. 11.1 outputs $\sigma$-soft inclusion correctly by virtue of Lemma 1.1. Next we prove that it also correctly claims exclusion.
@@
Now, if Alg. 11.1 outputs exclusion, then $1+\frac{d}{v_{g}}>\sigma^{h}$ for $g=0,1,\dots,q-1$, where $\sigma^{h}>1+2d\sqrt{q}$ according to initialization of Alg. 11.1. Hence $\frac{1}{v_{g}}>2\sqrt{q}$, and so the assumptions of Prop. 11.1 hold for $p_{h}(x)$ replacing $p(x)$. Therefore, the theorem implies that the disc $D(0,1)$ contains no zeros of $p_{h}(x)$, but then it contains no zeros of $p(x)$ as well (see Observation 10.2 (iii)), and so exclusion is correct.
@@
Alg. 11.1 is a $\sigma$-soft e/i test, certifying inclusion at the cost of evaluation of polynomial $p(x)$ at $2hj$ points $x$ or exclusion at the cost of evaluation at $2hq$ points $x$ for three integers $h$, $j$, and $q$ such that $1\leq j\leq q$, $q>m$, and $h:=\lceil\log_{\sigma}(1+2d\sqrt{q})\rceil$, $h=O(\log(d))$ for $\sigma-1$ exceeding a positive constant.
@@ 11.3 A faster deterministic e/i test
Prop. 10.2, for $p_{h}$ replacing $p$, and the assumed bound $\frac{d-m}{\theta^{h}-1}\leq\frac{1}{6\sqrt{q}}$ together imply
@@
Alg. 11.2 claims inclusion if $v_{g}\geq\frac{1}{3\sqrt{q}}$ for some $g$, but then
@@
Now Observation 10.2 (iii) implies that Alg. 11.2 claims inclusion correctly.
@@
Combine these bounds to obtain $\max_{g=0}^{q-1}\Big{|}\frac{f^{\prime}(\zeta_{q}^{g})}{f(\zeta_{q}^{g})}\Big{|}<\frac{1}{2\sqrt{q}}$. To complete the proof, apply Prop. 11.1 for $f(x)$ replacing $p(x)$ and for $m$ replacing $d$. ∎
Readily extend Prop. 11.2 to Alg. 11.2 by decreasing its bounds on $q$ and $h$ provided that #$\#(D(0,\theta))=m$ for $\theta\geq 1+6(d-m)\sqrt{q}$.
@@
Alg. 11.2 is a $\sigma$-soft e/i test: it certifies either inclusion or exclusion at the cost of evaluation of $p(x)$ at $2hj$ or $2hq$ points $x$, respectively, for three integers $h$, $j$, and $q$ such that $1\leq j\leq q$, $q>m:=\#(D(0,\theta)),~{}\sigma^{h}>1+6m\sqrt{q}$, $\theta\geq 1+6(d-m)\sqrt{q}$, and $h:=\lceil\log_{\sigma}(1+6m\sqrt{q})\rceil$, so that $h=O(\log(m))$ if $\sigma-1$ exceeds a positive constant.
@@ 11.4 A randomized e/i test
13.4.1. Introductory comments. Would Algs. 11.1 and 11.2 stay correct if we apply them at fewer evaluation points? They have consistently stayed correct for $q$ of order $\log^{2}(d)$ in extensive numerical tests of [57, 60] applied to $p(x)$ rather than to $p_{i}(x)$ for $i>0$ (that is, used no root-lifting). The tests reported exclusion where the values $|s_{h,q}|$ were reasonably small for $s_{h,q}:=\frac{1}{q}\sum_{g=0}^{q-1}\zeta^{(h+1)g}~{}\frac{p^{\prime}(\zeta^{g})}{p(\zeta^{g})},~{}{\rm for}~{}h=0,1,2$ (cf. Eqn. 3.2 in Part II), and reported inclusion otherwise. Clearly, $|s_{h,q}|\leq\delta$ for all $h$ if $|p(\zeta^{g})|\leq\delta/q$ for $g=0,1,\dots,q-1$, while inclusion is certified unless $|p(\zeta^{g})|$ is small for all $g$.
@@
Let $q=m-1$, $p(x)=(1-ux^{d-m})(x^{m}-mx+u)$, $u\approx 0$ but $u\neq 0$. (i) First let $m=d$. Then $p(x)=(1-u)(x^{d}-dx+u)$, $p^{\prime}(x)=(1-u)d(x^{d-1}-1)$ and NIR$(x)=0$ for $x$ being any $(d-1)$-st roots of unity. Extension of our e/i test to $q=d-1$ would imply that #$\#(D(0,1))=0$, while $p(y)=0$ for $u\approx 0$ and some $y\approx 0$. (ii) For $0<m<d$ notice that $p^{\prime}(x)=uf(x)+m(1-ux^{d-m})(x^{m-1}-1)$ where $f(x)=(d-m)x^{d-m-1}(x^{m}-mx+u)$ and so $uf(x)\approx 0$ while $1-ux^{d-m}\approx 1$ for $|x|\leq 1$ and $u\approx 0$. Therefore, $p^{\prime}(x)\approx 0$ where $x$ is any $(m-1)$-st root of unity, and then again extension of our e/i test to $q=m-1$ would imply that #$\#(D(0,1))=0$, while $p(y)=0$ for $u\approx 0$ and some $y\approx 0$.
Clearly, [57, 60] have never encountered worst case inputs in their extensive tests. Next we formally support this observation by introducing probability structure in the space of polynomials $p(x)$ and $t(x)=p(\frac{x-c}{\rho})$. Shifts destroy all customary probability distributions in the space of the coefficients of $p(x)$, but we prove correctness whp of e/i tests that evaluate $p(x)$ at $O(\log(d))$ under the Random Root Model.
@@
Fix any constant $\beta>1$ and a sufficiently large integer $h$ of order $\log(d)$. Then under the Random Root Model, Alg. 1.1(rrm) is correct with a probability at least $1-P$, where
@@
By virtue of Lemma 1.1 Alg. 1.1(rrm) outputs $\sigma$-soft inclusion correctly and therefore can be incorrect only where it outputs exclusion for a disc $D(0,1)$ containing a zero of $p$.
Let this zero be $z_{1}$, say. Then fix the values of the random variables $z_{2},\dots,z_{d}$ and $x_{1}\in C(0,1)$ and let $\gamma$ denote the area (Lebesgue’s measure) of the set $\mathbb{S}$ of the values of the variable $z_{1}$ for which
@@
Then Alg. 1.1(rrm) is incorrect with a probability at most
@@
Therefore $\lambda_{g+1}+\bar{z}^{g+1}=\lambda_{g}z+\bar{z}^{g}z$, and so
@@
For a fixed $\sigma>1$, e.g., for $\sigma=1.2$, choose a constant $\beta>1$ and sufficiently large $g=O(\log(d))$ such that $\gamma<\pi$ in (11.4). Then under the Random Root Model the probability $P$ of the output error for Alg. 1.1(rrm) is less than a constant $\nabla<1$. We can decrease the error probability below $\nabla^{v}$ by applying the algorithm $v$ times, in particular below $d^{a}$ for any fixed positive constant $a$ and for $v=O(\log(d))$.
Cor. 1.2 reduces classical subdivision root-finder to performing $O(mb)$ soft e/i tests, but in Part IV we decrease that bound to $w=O(m\log(b))$ tests. These tests magnify the error bounds by at most a factor of $w$. Since $m\leq d$, we can ensure correctness of the resulting e/i test whp provided that $\log(b)=o(d^{a-1})$. We arrive at the following
@@
Under Random Root Model one can perform all $O(m\log(b)))$ soft Las Vegas randomized e/i tests involved in a subdivision root-finder which ensure output error bound $\epsilon=1/2^{b}$ whp provided that $\log(b)=d^{O(1)}$ and that every test is reduced to $v=O(\log(d))$ applications of Alg. 1.1(rrm).
@@
It is tempting to combine root-lifting or root-squaring with algorithms of [111], which for approximate Cauchy integral (1.9), to obtain alternative randomized e/i test supporting Cor. 11.1, but for that we would have to raise disc isolation from order of $1+1/m$ to a constant, and this would require lifting of order $m$.
@@ 11.5 Counting the evaluation queries of our root-finders
Combine Cors. 1.2 and 11.1 and Prop. 11.4 and obtain
@@
We can reduce root-finding to deterministic evaluation of $p(x)$ at $O(m^{2}b\log(d))$ points as well as evaluation at $O(mb\log^{2}(d))$ points under the Random Root Model.
@@ 12 Reduction of $m$-test to e/i tests
Our algorithm in Sec. 12.1 reduces $m$-test to e/i tests motivated by application in Part IV. For these application our slower alternative reduction in Sec. 12.2 is inferior, but we include it because it relaxes isolation assumption and works for any disc, which can be of independent interest.

### v36

@@ 13.4 The distance to a closest root and fast inclusion tests
For a complex $c$ and positive integers $j$ and $h$, Eqn. (3.1) and Observation 13.2 combined imply the following relationships independent of $p_{d}$:
@@ 14.1 Overview
Map (1.3) and Cor. 3.1 together reduce an e/i test for any disc to the test for the unit disc $D=D(0,1)$, and we devise soft e/i tests for that disc assuming that #$\#(D(0,2))\leq m$ for a fixed $m$, $1\leq m\leq d$. Our Alg. 14.1 of Sec. 14.2 and Alg. 14.2 of Sec. 14.3 are deterministic specifications of Alg. 3.1, which is always correct when it claims inclusion, by virtue of Lemma 3.1. We prove that both of these specifications are also correct when they output exclusion. Namely, Algs. 14.1 and 14.2 evaluate ${\rm NIR}(x)$ at $hq$ equally spaced points of the unit circle $C(0,1)$ for $h=O(\log(d))$ and correctly claim exclusion if at all these points $|{\rm NIR}(x)|\leq\frac{1}{2\sqrt{q}}$ for $q>d$ or if $|{\rm NIR}(x)|\leq\frac{1}{3\sqrt{q}}$ for $q>m$, respectively. Our Example 14.1 shows that we cannot devise such deterministic algorithms for $q$ decreased below $m$, but in Sec. 14.4 we devise Las Vegas randomized e/i tests that evaluate polynomial $p(x)$ at $O(\log(d))$ points $x$ overall under the Random Root Model.
@@ 14.2 Basic deterministic e/i test
INPUT: a black box polynomial $p(x)$ of a degree $d$ given by an oracle for its evaluation and $\sigma$, $1<\sigma<\sqrt{2}$.
@@
Alg. 14.1 outputs $\sigma$-soft inclusion correctly by virtue of Lemma 3.1. Next we prove that it also correctly claims exclusion.
@@
Now, if Alg. 14.1 outputs exclusion, then $1+\frac{d}{v_{g}}>\sigma^{h}$ for $g=0,1,\dots,q-1$, where $\sigma^{h}>1+2d\sqrt{q}$ according to initialization of Alg. 14.1. Hence $\frac{1}{v_{g}}>2\sqrt{q}$, and so the assumptions of Prop. 14.1 hold for $p_{h}(x)$ replacing $p(x)$. Therefore, the theorem implies that the disc $D(0,1)$ contains no zeros of $p_{h}(x)$, but then it contains no zeros of $p(x)$ as well (see Observation 13.2 (iii)), and so exclusion is correct.
@@
Alg. 14.1 is a $\sigma$-soft e/i test, certifying inclusion at the cost of evaluation of polynomial $p(x)$ at $2hj$ points $x$ or exclusion at the cost of evaluation at $2hq$ points $x$ for three integers $h$, $j$, and $q$ such that $1\leq j\leq q$, $q>m$, and $h:=\lceil\log_{\sigma}(1+2d\sqrt{q})\rceil$, $h=O(\log(d))$ for $\sigma-1$ exceeding a positive constant.
@@ 14.3 A faster deterministic e/i test
Prop. 13.2, for $p_{h}$ replacing $p$, and the assumed bound $\frac{d-m}{\theta^{h}-1}\leq\frac{1}{6\sqrt{q}}$ together imply
@@
Alg. 14.2 claims inclusion if $v_{g}\geq\frac{1}{3\sqrt{q}}$ for some $g$, but then
@@
Now Observation 13.2 (iii) implies that Alg. 14.2 claims inclusion correctly.
@@
Combine these bounds to obtain $\max_{g=0}^{q-1}\Big{|}\frac{f^{\prime}(\zeta_{q}^{g})}{f(\zeta_{q}^{g})}\Big{|}<\frac{1}{2\sqrt{q}}$. To complete the proof, apply Prop. 14.1 for $f(x)$ replacing $p(x)$ and for $m$ replacing $d$. ∎
Readily extend Prop. 14.2 to Alg. 14.2 by decreasing its bounds on $q$ and $h$ provided that #$\#(D(0,\theta))=m$ for $\theta\geq 1+6(d-m)\sqrt{q}$.
@@
Alg. 14.2 is a $\sigma$-soft e/i test: it certifies either inclusion or exclusion at the cost of evaluation of $p(x)$ at $2hj$ or $2hq$ points $x$, respectively, for three integers $h$, $j$, and $q$ such that $1\leq j\leq q$, $q>m:=\#(D(0,\theta)),~{}\sigma^{h}>1+6m\sqrt{q}$, $\theta\geq 1+6(d-m)\sqrt{q}$, and $h:=\lceil\log_{\sigma}(1+6m\sqrt{q})\rceil$, so that $h=O(\log(m))$ if $\sigma-1$ exceeds a positive constant.
@@ 14.4 A randomized e/i test
14.4.1. Introductory comments. Would Algs. 14.1 and 14.2 stay correct if we apply them at fewer evaluation points? They have consistently stayed correct for $q$ of order $\log^{2}(d)$ in extensive numerical tests of [60, 63] applied to $p(x)$ rather than to $p_{i}(x)$ for $i>0$. The tests reported exclusion where the values $|s_{h,q}|$ were reasonably small for $s_{h,q}:=\frac{1}{q}\sum_{g=0}^{q-1}\zeta^{(h+1)g}~{}\frac{p^{\prime}(\zeta^{g})}{p(\zeta^{g})},~{}{\rm for}~{}h=0,1,2$ (cf. Eqn. 7.2 in Part II), and reported inclusion otherwise. Clearly, $|s_{h,q}|\leq\delta$ for all $h$ if $|p(\zeta^{g})|\leq\delta/q$ for $g=0,1,\dots,q-1$, while inclusion is certified unless $|p(\zeta^{g})|$ is small for all $g$.
@@
Let $q=m-1$, $p(x)=(1-ux^{d-m})(x^{m}-mx+u)$, $u\approx 0$ but $u\neq 0$. (i) First let $m=d$. Then $p(x)=(1-u)(x^{d}-dx+u)$, $p^{\prime}(x)=(1-u)d(x^{d-1}-1)$ and NIR$(x)=0$ for $x$ being any $(d-1)$-st roots of unity. Then our e/i test for $q=d-1$ would imply that #$\#(D(0,1))=0$, while $p(y)=0$ for $u\approx 0$ and some $y\approx 0$. (ii) For $0<m<d$ notice that $p^{\prime}(x)=uf(x)+m(1-ux^{d-m})(x^{m-1}-1)$ where $f(x)=(d-m)x^{d-m-1}(x^{m}-mx+u)$ and so $uf(x)\approx 0$ while $1-ux^{d-m}\approx 1$ for $|x|\leq 1$ and $u\approx 0$. Therefore, $p^{\prime}(x)\approx 0$ where $x$ is any $(m-1)$-st root of unity, and then again extension of our e/i test to $q=m-1$ would imply that #$\#(D(0,1))=0$, while $p(y)=0$ for $u\approx 0$ and some $y\approx 0$.
Clearly, [60, 63] have never encountered worst case inputs in their extensive tests. Next we formally support this observation by introducing probability structure in the space of polynomials $p(x)$ and $t(x)=p(\frac{x-c}{\rho})$. Shifts destroy all customary probability distributions in the space of the coefficients of $p(x)$, but we prove that our e/i tests that evaluate $p(x)$ at $O(\log(d)$ points are correct whp under the Random Root Model.
@@
Fix any constant $\beta>1$ and a sufficiently large integer $h$ of order $\log(d)$. Then under the Random Root Model, Alg. 3.1(rrm) is correct with a probability at least $1-P$, where
@@
By virtue of Lemma 3.1 Alg. 3.1(rrm) outputs $\sigma$-soft inclusion correctly and, therefore, can be incorrect only where it outputs exclusion for a disc $D(0,1)$ containing a root.
Let this root be $z_{1}$, say. Then fix the values of the random variables $z_{2},\dots,z_{d}$ and $x_{1}\in C(0,1)$ and let $\gamma$ denote the area (Lebesgue’s measure) of the set $\mathbb{S}$ of the values of the variable $z_{1}$ for which
@@
Then Alg. 3.1(rrm) is incorrect with a probability at most
@@
Therefore, $\lambda_{g+1}+\bar{z}^{g+1}=\lambda_{g}z+\bar{z}^{g}z$, and so
@@
For a fixed $\sigma>1$, e.g., for $\sigma=1.2$, choose a constant $\beta>1$ and sufficiently large $g=O(\log(d))$ such that $\gamma<\pi$ in (14.4). Then under the Random Root Model the probability $P$ of the output error for Alg. 3.1(rrm) is less than a constant $\nabla<1$. We can decrease the error probability below $\nabla^{v}$ by applying the algorithm $v$ times, in particular below $d^{a}$ for any fixed positive constant $a$ and for $v=O(\log(d))$.
Based on Cor. 2.2 we reduce classical subdivision root-finder to performing $O(mb)$ soft e/i tests, but in Part IV we decrease that bound to $w=O(m\log(b))$ tests. These tests magnify the error bounds by at most a factor of $w$. Since $m\leq d$, we can ensure correctness of the resulting e/i test whp provided that $\log(b)=o(d^{a-1})$. We arrive at the following
@@
Under Random Root Model one only need to apply a subdivision root-finder that involves $O(m\log(b)))$ soft Las Vegas randomized e/i tests to ensure the output error bound $\epsilon=1/2^{b}$ whp provided that $\log(b)=d^{O(1)}$ and that every test is reduced to $v=O(\log(d))$ applications of Alg. 3.1(rrm).
@@
It is tempting to combine root-lifting or root-squaring with algorithms of [115], which approximate Cauchy integral (3.4), to obtain alternative randomized e/i test supporting Cor. 14.1, but for that we would have to raise disc isolation from order of $1+1/m$ to a constant, and this would require too costly lifting of order $m$.
@@ 14.5 Counting the evaluation queries of our root-finders
Combine Cors. 2.2 and 14.1 and Prop. 14.4 and obtain
@@
We can reduce root-finding to deterministic evaluation of $p(x)$ at $O(m^{2}b\log(d))$ points as well as to evaluation at $O(mb\log^{2}(d))$ points under the Random Root Model.
@@ 15 Reduction of $m$-test to e/i tests
Our two algorithms of Secs. 15.1 and 15.2 reduce $m$-test for a disc $D$ to bounded number of e/i tests. Namely, the algorithm of Sec. 15.2 uses just a single e/i test where $m=d$ but up to order of $\log(d)$ e/i tests otherwise, while the algorithm of Sec. 15.1 uses $O(1)$ e/i tests for any $m$.
15.1 $m$-test by means of e/i tests for any disc
Our alternative reduction of $m$-test for $1\leq m\leq d$ to a bounded number of e/i tests can be applied in the case of any disc, not necessarily isolated. If at least one of the e/i tests fails, then definitely there is a root in a narrow annulus about the boundary circle of an input disc, and hence the internal disc of the annulus would not contain $m$ roots. Otherwise we would obtain a root-free concentric annulus about the boundary circle of an input disc and then reduce $m$-test to root-counting in an isolated disc, which we can perform at a low cost (see, e.g., claim (ii) of Prop. 7.1).
Algorithm 15.1.
Reduction of $m$-test to e/i tests.
INPUT: a polynomial $p(x)$ of a degree $d$, two values $\bar{\theta}>1$ and $\bar{\sigma}>1$, an integer $m$ such that $1\leq m\leq d$ and #$\#(D(0,\bar{\sigma}))\leq m$, and two algorithms $\mathbb{ALG}~{}0$ and $\mathbb{ALG}~{}1$, which for a black box polynomial perform $\bar{\sigma}$-soft e/i test for in any disc and root-counting in an isolated disc, respectively.
INITIALIZATION (see Fig. 6): Fix a positive $\rho$ such that $\bar{\sigma}=O(\rho)$, say, $\rho=\bar{\sigma}/5$, and compute the integer
$v=\lceil 2\pi/\phi\rceil$ (15.1)
where $\phi$ is the angle defined by an arc of $C(0,1)$ with end points at the distance $\rho$, and so $v=O(1)$ for $\bar{\sigma}=O(\rho)$. Then compute the values
$c_{j}=\exp(\phi_{j}{\bf i}),~{}j=0,1,\dots,v,$ (15.2)
$c_{1}$$c_{2}$$c_{3}$$c_{4}$$c_{5}$$c_{6}$$c_{7}$$c_{8}$$c_{9}$$c_{0}$
Figure 6: $v=9$, $|c_{j}-c_{j-1}|=\rho>|c_{9}-c_{0}|$ for $j=1,2,\dots,9$.
To prove correctness of Alg. 15.1 we need the following result.
Theorem 15.1.
Proof.
$c_{0}$$c_{1}$$z$$z^{\prime}$
Corollary 15.1.
Correctness of Alg. 15.1. (i) If algorithm $\mathbb{ALG}~{}1$ outputs inclusion in at least one of its applications, then one of the discs $D(c_{j},\bar{\sigma}\rho)$ contains a root, and then Alg. 15.1 correctly reports that $r_{d-m+1}\geq 1-\sigma\bar{\rho}$.
(ii) Otherwise Cor. 15.1 holds, and then correctness of Alg. 15.1 follows from correctness of $\mathbb{ALG}~{}1$.
Computational complexity. Alg. 15.1 invokes algorithm $\mathbb{ALG}~{}0$ at most $v+1$ times for $v=O(1)$ of (15.1). If $\mathbb{ALG}~{}0$ outputs exclusion $v+1$ times, then algorithm $\mathbb{ALG}~{}1$ is invoked and runs at a dominated cost.
Remark 15.1.
One can modify Alg. 15.1 as follows (see [63]): instead of fixing a positive $\rho$ and then computing $v$ of (15.1) fix an integer $v>0$, apply algorithm $\mathbb{ALG}~{}0$ at $v$ equally spaced points $c_{j}=\exp(\frac{2j\pi{\bf i}}{v}),~{}j=0,1,\dots,v-1$, let $\phi=\frac{2\pi}{v}$, and then define $\rho$. We also arrive at this algorithm if we compute or guess this value $\rho$ and then apply Alg. 15.1, but notice that $c_{v}=c_{0}$ for $c_{j}$ of (15.2) and only invoke algorithm $\mathbb{ALG}~{}0$ for $c_{0},\dots,c_{v-1}$.